# Improving Temporal Understanding Logic Consistency in Video-Language Models via Attention Enhancement

### v1

@@ 4.2.3 Discussion
1). The improvements brought by TCAS are evident across different models, training datasets, and benchmarks. As shown in Table 3, TCAS achieves favorable results in all comparisons, indicating that TCAS is not dependent on specific models, benchmarks, or training annotations and can robustly enhance model performance and consistency. The performance improvement on ActivityNet-CON is relatively smaller, which is due to the noisier data distribution of this dataset, where event descriptions are presented in long sentences that deviate from the general understanding of the concept of events. Consequently, all methods and models perform worse on this dataset, making the smaller improvement acceptable.
2). TCAS improves both grounding and consistency performance. On the classic VTG benchmarks, Charades-STA and ActivityNet-Captions, TCAS achieves SOTA performance among Video-LLMs methods. This suggests that inconsistent temporal understanding is a key limitation for the grounding performance of these models. By enhancing attention discriminability, TCAS improves temporal logic consistency, thereby boosting grounding performance and advancing video temporal understanding.
@@ 4.3 Ablation Study
The impact of different hyperparameters on model performance varies. The parameters $t$ and thrthr primarily determine the adjustment scope of TCAS on the aspect of heads and tokens, respectively, while $m$ and $w_{ae}$ control the intensity of TCAS enhancement. As shown in Table 5, increasing $t$ leads to a significant drop in G. and R-G. subsets, while a larger thrthr results in a substantial decline in S-G., H-V., and C-V. subsets, indicating that these two parameters are most critical for model performance. Overall, our method is more sensitive to scope-related hyperparameters while maintaining robustness to intensity-related hyperparameters.
@@ 4.4 Case Study
To visually demonstrate the effectiveness of our method, we select several prediction results from Charades-CON Jung et al. (2024) for a case study. For the example shown in Figure 6, our model successfully grounds query events in the video and correctly answers the questions, demonstrating improved temporal reasoning capabilities compared to baseline methods. This result highlights the potential of our approach to enhance the understanding of temporal dynamics in video content. More examples are provided in the Appendix B.5.
4.5 Statistical Analysis
To confirm that TCAS improves consistency by enhancing attention discriminability, rather than through unintended artifacts, we analyze the distribution of the top-32 attention discriminability scores for both the baseline and TCAS-enhanced models on the Charades-CON dataset. As depicted in the histogram in Figure 7, the statistical analysis reveals a distinct rightward shift in the score distribution after enhancement. This demonstrates that TCAS effectively boosts attention discriminability, which consequently improves consistency, thereby reinforcing the validity of our conclusions.
@@ 4.5 Statistical Analysis
5 Related Work
5.1 Video-LLMs on VTG
Recent Video-LLMs excel in fine-grained temporal understanding tasks like VTG by independently encoding timestamps and interacting with video information to capture precise temporal segments, marking a significant improvement over traditional VTG methods by enabling diverse natural language outputs Guo et al. (2025); Ren et al. (2024); Hu et al. (2024). New benchmarks for fine-grained temporal understanding have been proposed to analyze their limitations Fu et al. (2024); Jung et al. (2024). For example, the consistency benchmark by Jung et al. Jung et al. (2024) highlights response inconsistency. Unlike prior work, we focus on the internal representations of Video-LLMs rather than solely developing new benchmarks for evaluation, providing deeper insights into their temporal cognition.
5.2 Interpretability Analysis in MLLMs
In recent years, interpretability of LLMs has emerged as a key research focus to enhance model transparency and controllability Rai et al. (2025). However, the understanding of MLLMs remains limited Lin et al. (2025). Nikankin Nikankin et al. (2025) and Venhoff Venhoff et al. (2025) investigated MLLMs’ image understanding using circuit and SAE, respectively. Li et al. Li et al. (2025) probed visual features and identified the LLM decoder as the bottleneck for temporal reasoning, but did not explore the root causes of this limitation. From a consistency perspective, we offer a deeper explanation, attributing the decoder’s limitations to the insufficient discriminability of cross-modal attention heads.
6 Conclusion
In this paper, we first analyze the role of cross-modal attention heads in establishing temporal cognition and then validate the causal relationship between attention discriminability and temporal logic consistency through correlation analysis and causal intervention.
Based on these insights, we introduce a novel method, TCAS, that optimizes attention distributions to improve temporal discriminability. Our experimental results demonstrate that our method effectively enhances temporal logic consistency in Video-LLMs, achieving SOTA grounding performance and consistency performance across various settings, while other extensive experiments further demonstrate its reliability.
@@ A.3 Details of Analysis on EOJ Task
Next, we visualize the attention patterns of key heads on several samples, as shown in Figure 12. We find that the attention patterns of these heads are consistent with the analysis in the main text, establishing a mapping relationship between the query tokens of event queries and the video key tokens corresponding to the time range of events through attention scores, thus achieving cross-modal temporal relationship alignment.
Figure 12: Attention patterns of key heads in the EOJ task.
A.3.2 Correlation Analysis.
@@
$c_{v}=\frac{\sum_{q\in Q_{v}}F1(r_{v,q})}{|Q_{v}|},\quad v\in\mathcal{D}$ (4)
where $c_{v}$ indicates the model’s Consistency Score for video sample $v$, $r_{v,q}$ denotes the model’s response to question $q$ for video sample $v$, $Q_{v}$ refers to the set of questions associated with video sample $v$, and $\mathcal{D}$ represents the dataset of all video samples.
The Attention Discriminability Score $S^{h,v}_{disc.}$ is defined as the Kullback-Leibler divergence Kullback and Leibler (1951) between the average attention score distributions of two events, specifically:
$P^{h,v}_{e}=\frac{\sum_{q\in T_{e}}A^{h,v}_{q,V}}{\sum_{k\in V}\sum_{q\in T_{e}}A^{h,v}_{q,k}}$ (5)
$S^{h,v}_{disc.}=P^{h,v}_{e_{1}}\log\frac{P^{h,v}_{e_{1}}}{P^{h,v}_{e_{2}}}+P^{h,v}_{e_{2}}\log\frac{P^{h,v}_{e_{2}}}{P^{h,v}_{e_{1}}}$ (6)
$S^{v}_{disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{disc.}$ (7)
where $P^{h,v}_{e}$ denotes the average attention score distribution of all text tokens of event $e$ to the set of all visual tokens $V$, $T_{e}$ represents the set of text tokens of event $e$, $A^{h,v}_{q,V}$ is the attention score from text token as query token $q$ to all visual tokens $V$ as key tokens, $S^{h,v}_{disc.}$ is the temporal discriminability score of attention head $h$ for video sample $v$ containing two events $e_{1}$ and $e_{2}$, $H_{t}$ means the set of the top $t$ cross-modal score attention heads, and $S^{v}_{disc.}$ is the average temporal discriminability score of the top $t$ cross-modal attention heads for video sample $v$.
The correlation analysis of attention discriminability scores and consistency scores is shown in Figure 13 as a scatter plot with the least-squares regression line. We observe that as the consistency score increases, the attention discriminability score also increases significantly, indicating a clear positive correlation between attention discriminability and consistency performance. The corresponding violin plot is shown in Figure 5 of the main text.
@@
Appendix B More Details about Experiments
B.1 Backbone Models Introduction
Qwen2.5-VL Bai et al. (2025) is a representative Video-LLMs featuring a standard cross-modal architecture. We select Qwen2.5-VL to further validate our findings from the previous section. Video-LLaMA Zhang et al. (2023) is a classic video-LLM using Qformer and video Qformer for video-text alignment. TimeChat Ren et al. (2024), a SOTA video-LLM for the VTG task, integrates timestamp information into its architecture, enhancing its ability to understand the temporal dynamics of video content. We include Video-LLaMA and TimeChat to ensure a fair comparison with the current SOTA method.
B.2 Datasets Introduction and Examples
TimeIT
TimeIT is a dataset proposed by Ren et al. Ren et al. (2024) for the various temporal understanding tasks including description, video grounding and temporal reasoning. The splits Charades-VTG and ActivityNet-VTG of it is used in our experiments. These two splits contain 9,848 video clips with 12,408 question-answer pairs and 14,420 video clips with 35,692 question-answer pairs respectively.
VTune
VTune is a dataset reported as a method in the original paper Jung et al. (2024). By constructing event verification queries and temporal verification queries annotation data, it effectively enhances the consistency of Video-LLMs from a view of synthetic data augmentation. Two splits Charades-VTune and ActivityNet-VTune contains 9,848 video clips with 99,244 question-answer pairs and 14,420 video clips with 205,510 question-answer pairs.
CON
Charades-CON and ActivityNet-CON are two benchmarks proposed by Jung et al. Jung et al. (2024) to evaluate the consistency of Video-LLMs in temporal understanding tasks. These benchmarks contain 500 video clips with 707 question-answer pairs and 1,422 question-answer pairs.
B.3 Evaluation Metrics
Following the previous works Jung et al. (2024), we use these metrics to evaluate the performance of our method:
  • mIoU: The mean IoU of all questions in the dataset.
B.4 Other Baselines
We also compare our method with the following closed-source models: GPT-4o OpenAI (2024) Gemini-1.5 Flash Team (2024) in Table 6. It is worth noting that these results are reported in Jung et al. (2024) and are listed here for reference only.
@@ G R-Ground S-Ground H-Verify C-Verify
Table 6: Results of other models without tuning, listed for reference only Jung et al. (2024).
@@ B.5 Case Study
In this section, we present more prediction examples in Figure 14 to illustrate the effectiveness of our method in enhancing temporal consistency in Video-LLMs. Our model successfully grounds key events and answers questions about the video, demonstrating improved temporal capabilities and consistency compared to baseline models.

### v2

@@ 5.2.3 Discussion
1). The improvements brought by TCAS are evident across different models, training datasets, and benchmarks. As shown in Table 3, TCAS achieves favorable results in all comparisons, indicating that TCAS is not dependent on specific models, benchmarks, or training annotations and can robustly enhance model consistency. The performance improvement on ActivityNet-CON is relatively smaller, which is due to the noisier data distribution of this dataset, where event descriptions are presented in long sentences that deviate from the general understanding of the concept of events. Consequently, all methods and models perform worse on this dataset, making the smaller improvement acceptable.
2). TCAS not only improves consistency as anticipated but also yields unexpected gains in grounding performance. This suggests that inconsistent temporal understanding is a potential factor limiting the grounding of Video-LLMs. We attribute this improvement to the regularization effect of TCAS. While standard SFT often fits dataset biases to satisfy token-level objectives—relying on shortcuts like language priors [31, 30] rather than true grounding evidence—TCAS effectively suppresses such non-essential cues. By forcing the model to utilize discriminative attention on temporal tokens for reasoning, our method highlights authentic temporal evidence and improves generalization, ultimately advancing the model’s fundamental temporal comprehension.
@@ 5.3 Ablation Study
The impact of different hyperparameters on model performance varies. The parameters $t$ and thrthr primarily determine the adjustment scope of TCAS on the aspect of heads and tokens, respectively, while $m$ and $w_{ae}$ control the intensity of TCAS enhancement. As shown in Table 5, increasing $t$ leads to a significant drop in G. and R-G. subsets, while a larger thrthr results in a substantial decline in S-G., H-V., and C-V. subsets, indicating that these two parameters are most critical for model performance. Overall, our method is more sensitive to scope-related hyperparameters while maintaining robustness to intensity hyperparameters.
@@ 5.4 Case Study
To visually demonstrate the effectiveness of our method, we select several prediction results from Charades-CON [9] for a case study. For the example shown in Figure 6, our model successfully grounds query events in the video and correctly answers the questions, demonstrating improved temporal reasoning capabilities compared to baseline methods. This result highlights the potential of our approach to enhance the understanding of temporal dynamics in video content. More examples are provided in the Appendix B.5.
5.5 Robustness to Video Duration
We evaluate the robustness of TCAS across different video durations on Charades-CON. As shown in Table 6, while the baseline performance typically declines in longer videos due to increased temporal complexity, TCAS yields progressively larger gains. Specifically, for videos exceeding 40 seconds, TCAS achieves substantial improvements of +17.7 points and +14.8 points in Grounding (G.) and Rephrase-Grounding (R-G.), respectively. This highlights its efficacy in maintaining discriminative attention and temporal consistency in challenging, long-form video understanding.
Table 6: Robustness analysis across video durations on Charades-CON. $n$ is the sample count. Gains over SFT are shown in parentheses.
0–20s, n58 20–30s, n189 30–40s, n409 $\mathbf{>}$40s, n51
SFT TCAS SFT TCAS SFT TCAS SFT TCAS
G. 89.7 93.1+3.4 78.8 84.1+5.3 74.1 82.4+8.3 58.8 76.5+17.7
R-G. 84.5 86.2+1.7 72.3 76.4+4.1 66.3 73.9+7.6 51.2 66.0+14.8
S-G. 46.6 55.2+8.6 35.5 38.1+2.6 34.1 39.7+5.6 20.9 25.5+4.6
5.6 Visualization Analysis
We analyzed the top-32 attention discriminability scores to confirm that TCAS’s consistency improvements stem from enhanced discriminability, rather than unintended artifacts. Figure 7 shows a clear rightward shift in the score distribution for the TCAS-enhanced model compared to the baseline. This result demonstrates that TCAS directly boosts attention discriminability, which consequently improves consistency, thus validating our approach, thereby reinforcing the validity of our conclusions.
@@ 5.6 Visualization Analysis
6 Conclusion and Limitation
In this paper, we identify and validate a causal link between cross-modal attention discriminability and temporal logic consistency in Video-LLMs through correlation analysis and targeted intervention. Motivated by this finding, we introduce TCAS, a method that optimizes attention distributions to enhance temporal discriminability. Experiments demonstrate that TCAS significantly improves temporal logic consistency across various models. We acknowledge that our focus on logical inconsistency may not encompass all facets of temporal understanding.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. U21B2009) and the Foundation Program of the Key Laboratory of Science and Technology on Complex Electronic System Simulation (No. 614201001032203).
Supplementary Material
In this file, we provide additional details as follows:
  1. A.
    More Details about Analysis
    1. 1.
      More Attention Patterns. This section presents additional attention patterns of key attention heads in TimeChat, further supporting the analysis in Section 2.1 of the main text.
    2. 2.
      Correlation Analysis Details. This section provides the fitting curves and complete distributions of the correlation analysis, as well as further analysis.
    3. 3.
      Details of Analysis on EOJ Task. This section provides additional details on the analysis of the Event Order Judgement (EOJ) task, including the methodology, implementation details, results, and discussion.
  2. B.
    More Details about Experiments
    1. 1.
      Backbone Models Introduction. This section provides a detailed introduction to the backbone models used in our experiments, along with the reasons for their selection.
    2. 2.
      Datasets Introduction and Examples. This section describes the datasets used in our experiments, including examples to illustrate their content and structure.
    3. 3.
      Evaluation Metrics. This section details the evaluation metrics used to assess the performance.
    4. 4.
      Other Baselines and Backbones. This section discusses additional baselines and backbone models that are considered in our experiments.
    5. 5.
      Case Study. This section presents some case studies to illustrate the effectiveness of our method in enhancing temporal consistency in Video-LLMs.
    6. 6.
      Further Ablation Studies on Sensitivity. This section provides additional ablation studies to further validate the components of our proposed method.
  3. C.
    More Discussions
    1. 1.
      Method Design Related Work. This section reviews related work on contrastive learning relevant to our method design.
    2. 2.
      Limitations and Future Work. This section discusses the limitations of our current work and outlines potential directions for future research.
@@ A.3.1 Detection and Interpretation
Next, we visualize the attention patterns of key heads on several samples, as shown in Figure 5 in main text. We find that the attention patterns of these heads are consistent with the analysis in the main text, establishing a mapping relationship between the query tokens of event queries and the video key tokens corresponding to the time range of events through attention scores, thus achieving cross-modal temporal relationship alignment.
A.3.2 Correlation Analysis
@@
$c_{v}=\frac{\sum_{q\in Q_{v}}F1(r_{v,q})}{|Q_{v}|},\qquad v\in\mathcal{D}$ (5)
where $c_{v}$ indicates the model’s EOJ Consistency Score for video sample $v$, $r_{v,q}$ denotes the model’s response to question $q$ for video sample $v$, $Q_{v}$ refers to the set of questions associated with video sample $v$, and $\mathcal{D}$ represents the dataset of all video samples.
Considering that the EOJ task involves multiple events, the EOJ Attention Discriminability Score is defined as the Kullback-Leibler divergence [11] between the average attention score distributions of two events as follows:
$P^{h,v}_{e}=\frac{\sum_{q\in T_{e}}A^{h,v}_{q,V}}{\sum_{k\in V}\sum_{q\in T_{e}}A^{h,v}_{q,k}}\text{,}$ (6)
$S^{h,v}_{eoj,disc.}=P^{h,v}_{e_{1}}\log\frac{P^{h,v}_{e_{1}}}{P^{h,v}_{e_{2}}}+P^{h,v}_{e_{2}}\log\frac{P^{h,v}_{e_{2}}}{P^{h,v}_{e_{1}}}\text{,}$ (7)
$S^{v}_{eoj,disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{eoj,disc.}\text{,}$ (8)
where $P^{h,v}_{e}$ denotes the average attention score distribution of all text tokens of event $e$ to the set of all visual tokens $V$, $T_{e}$ represents the set of text tokens of event $e$, $A^{h,v}_{q,V}$ is the attention score from text token as query token $q$ to all visual tokens $V$ as key tokens, $S^{h,v}_{eoj,disc.}$ is the EOJ temporal discriminability score of attention head $h$ for video sample $v$ containing two events $e_{1}$ and $e_{2}$, $H_{t}$ means the set of the top $t$ cross-modal score attention heads, and $S^{v}_{eoj,disc.}$ is the average eoj temporal discriminability score of the top $t$ cross-modal attention heads for video sample $v$.
The correlation analysis of attention discriminability scores and consistency scores is shown in Figure 13 as a scatter plot with the least-squares regression line. We observe that as the consistency score increases, the attention discriminability score also increases significantly, indicating a clear positive correlation between attention discriminability and consistency performance. The corresponding violin plot is shown in Figure 12.
Figure 12: Violin plot of analysis results on EOJ task.
@@ G R-Ground S-Ground H-Verify C-Verify
Table 7: Results of other models without tuning, listed for reference only [9].
Appendix B More Details about Experiments
B.1 Backbone Models Introduction
Qwen2.5-VL [1] is a representative Video-LLMs featuring a standard cross-modal architecture. We select Qwen2.5-VL to further validate our findings from the previous section. Video-LLaMA [28] is a classic video-LLM using Qformer and video Qformer for video-text alignment. TimeChat [19], a SOTA video-LLM for the VTG task, integrates timestamp information into its architecture, enhancing its ability to understand the temporal dynamics of video content. We include Video-LLaMA and TimeChat to ensure a fair comparison with the current SOTA method.
B.2 Datasets Introduction and Examples
TimeIT
TimeIT is a dataset proposed by Ren et al. [19] for the various temporal understanding tasks including description, video grounding and temporal reasoning. The splits Charades-VTG and ActivityNet-VTG of it is used in our experiments. These two splits contain 9,848 video clips with 12,408 question-answer pairs and 14,420 video clips with 35,692 question-answer pairs respectively.
VTune
VTune is a dataset reported as a method in the original paper [9]. By constructing event verification queries and temporal verification queries annotation data, it effectively enhances the consistency of Video-LLMs from a view of synthetic data augmentation. Two splits Charades-VTune and ActivityNet-VTune contains 9,848 video clips with 99,244 question-answer pairs and 14,420 video clips with 205,510 question-answer pairs.
CON
Charades-CON and ActivityNet-CON are two benchmarks proposed by Jung et al. [9] to evaluate the consistency of Video-LLMs in temporal understanding tasks. These benchmarks contain 500 video clips with 707 question-answer pairs and 1,422 question-answer pairs.
B.3 Evaluation Metrics
B.3.1 Evaluation Metrics Details
Following the previous works [9], we use these metrics to evaluate the performance of our method:
  • mIoU: The mean IoU of all questions in the dataset.
B.3.2 Metrics Rationality
For analysis in Section 3, we use a sample-wise consistency score (IoU product) for instance-level correlation. In our main experiments (Sec. 5), we follow Jung et al.’s protocol [9], reporting R@0.5, Acc, and their product with the original grounding score as Appendix Section B.3.1, rather than sample-wise score for strictly fair comparison with baselines. For intervention, we report per-setting IoU (Ori/R./S.) to isolate subtask effects.
The result using sample-wise consistency score (the analysis metric) is shown in Table 8. A simple intervention on a small number of attentional heads can slightly improve the consistency score, while low-intensity random intervention leads to a rapid decline in performance. Gains may look limited because most components are frozen and cannot co-adapt to the intervention, constraining immediate gains. TCAS mitigates this by retraining, allowing the improved attention discriminability to better translate into performance, further confirming the effectiveness of our method TCAS in enhancing temporal consistency.
Table 8: Results of sample-wise consistency score defined as the product of IoU and original grounding score. The $\alpha$ means the intervention intensity. ”Random Int.” and ”Ground Truth Int.” mean random intervention and intervention based ground-truth.
SFT w/o Int. Random Int. Ground Truth Int. TCAS
$\alpha$ 0.05 0.1 0.15 0.2 0.4 0.6 0.8 1 -
$c_{rg.}$ 65.4 39.6 2.1 0 66.0 65.4 64.1 62.9 61.2 70.0
$c_{sg.}$ 33.1 17.2 1.8 0 33.5 33.7 34.0 34.5 34.7 37.5
B.4 Results of Other Baselines and Backbones
We also compare our method with the following closed-source models: GPT-4o [16] Gemini-1.5 Flash [21] in table 7. It is worth noting that these results are reported in [9] and are listed here for reference only.
Furthermore, we conducted comparative experiments on the general VTG task using Qwen2.5-VL and Video-LLaMA, with the results presented in Table 9. In contrast to the improvements seen with TimeChat, our method does not consistently enhance performance on the general VTG task for these models. However, the general localization performance remains stable or slightly improves with the enhanced consistency, which we consider acceptable, as the primary motivation of our work is to improve consistency in an interpretable manner.
Table 9: More comparison results for VTG performance on Charades-STA and ActivityNet-Caption.
Method Charades-STA ActivityNet-Cap.
R@1,0.5 R@1,0.7 R@1,0.5 R@1,0.7
videollama (w. SFT) 37.1 20.1 34.3 19.1
videollama (w. TCAS) 37.5 20.1 35.2 20.0
Qwen2.5vl (w. SFT) 27.3 14.1 16.5 7.8
Qwen2.5vl (w. TCAS) 32.4 16.6 16.5 7.9
@@ B.5 Case Study
In this section, we present more prediction examples to illustrate the effectiveness of our method in enhancing temporal consistency in Video-LLMs. Our model successfully grounds key events and answers questions about the video, demonstrating improved temporal capabilities and consistency compared to baseline models.
B.6 Further Ablation Studies on Sensitivity
Figure 16: Performance sensitivity analysis of hyperparameters $t$ (top subplot) and thrthr (bottom subplot).
We conducted a more fine-grained sensitivity analysis for the relatively sensitive hyperparameters $t$ and thrthr. The performance variation with respect to these hyperparameters is shown in Figure 16. From the figure, we observe that: 1) As $t$ increases, both consistency scores exhibit an overall trend of first increasing and then decreasing, emphasizing the need for precise regulation of attention heads, as excessive intervention on heads weakly related to temporal discriminability can have negative effects. 2) When thrthr is small, both Rephrase Consistency Score and Shift Consistency Score are relatively high; however, when thrthr exceeds 0.1, performance rapidly declines. Because larger values of thrthr result in too few tokens being adjusted, which fails to effectively enhance attention discriminability.